Not yet answered

Points out of 2

Flag question

Question text

11.3 grams of ammonium chloride (NH4Cl) are dissolved in enough water to make 155 mL of solution. What is the resulting Molarity of the aqeuous solution? Give your answer to the correct number of significant figures. Type in only the numbers. Do not include the molarity unit (M or moles/L). Hint: you need to solve for the molar mass of ammonium chloride to do this calculation.

Answer:

Question 2

Not yet answered

Points out of 2

Flag question

Question text

10.0 mL of the aqeuous ammonium chloride solution from question 1 above is diluted by adding 15.0 mL of distilled water. What is the new Molarity of this diluted solution? Give your answer to the correct number of significant figures. Type in only the numbers. Do not include the molarity unit (M or moles/L).

Note: the answer will be a molarity less than 1. Make sure to type in the leading zero. For example, if your answer is 0.111, type in 0.111 with the "zero point", and NOT .111

Answer:

Question 3

Not yet answered

Points out of 3

Flag question

Question text

Water is a very polar liquid, which does not mix with toluene, a second liquid. Instead, when the two are combined, toluene floats on top of water.

Toluene is Choose...nonpolarpolar and Choose...less densemore dense than water.

A chromatography experiment is performed where a specific ink is dotted on filter paper, and a toluene mobile phase is used. The ink dot does not move with the toluene as the liquid moves through the paper. Instead the ink dot remains in its original position on the filter paper.

The ink is Choose...polarnonpolar .
Answer #1

1.

11.3 grams of ammonium chloride is dissolved in 155 mL of solution.

We have the molar mass of ammonium chloride = 53.5 g/mol

We can calculate the molarity of a solution using the formula

Molarity = (Given weight/ Molar mass) (1000/ Volume of solution in mL)

Thus the required molarity of ammonium chloride

= (11.3 g / 53.5 g/mol) (1000 mL / 155 mL) = 11300 / 8292.5 = 1.3627 M

Therefore the molarity of ammonium chloride is 1.36